Given the graph below, find MN. Round to

the nearest hundredth.


1
Expert's answer
2021-10-27T14:10:50-0400

The distance Formula: "d =\sqrt{(x2 - x1)\\^2 + (y2 - y1)\\^2 }"

First point M: "(-1,2)"

Second Point N: "(4,0)"

"MN =\sqrt{(4 - (-1))\\^2 + (0 - 2)\\^2 }\\
MN=\sqrt{(5)\\^2 + (-2)\\^2 }\\
MN=\sqrt{25+4}\\
MN=\sqrt{29}\\
MN= 5.39 (nearest\ hundredths)"